OpenStudy (anonymous):

The value of \[2+\sqrt{2+\sqrt{2+\sqrt{2.......}}}\]

OpenStudy (anonymous):

gimmick is to put \[x=2+\sqrt{2+\sqrt{2+\sqrt{2.......}}}\] \[x-2=\sqrt{2+\sqrt{2+\sqrt{2.......}}}\] \[(x-2)^2=2+\sqrt{2+\sqrt{2+\sqrt{2.......}}}=x\]

OpenStudy (anonymous):

that is solve \[(x-2)^2=x\] hard part is proving the limit exists,

OpenStudy (anonymous):

I've seen this one before. I think I found it using brute force last time, then someone said something like Limit = 2 + sqrt(Limit), or something like that.